# Transformer Layers as Painters

### v2

@@ 1 Introduction
To answer these questions we perform a series of experiments on pretrained LLMs. These include experimenting with variations on the standard transformer execution strategy, and measuring the impact of these variations on the models’ performance across a variety of benchmarks for both decoder-only (Llama) and encoder-only (BERT) models. Note that our experiments never involve finetuning or otherwise adjusting the models’ parameters (with the caveat that the GLUE evaluation standard procedure includes a finetuning step for our BERT-Large model) [^1] Our code is available at https://anonymous.4open.science/r/transformer-as-painter-CF60.
@@ 3.1 Do layers “speak the same language”?
Figure 3: Avg. cosine similarity between the hidden states of all 32 layers of Llama2-7B (Left) and all 40 layers of Llama2-13B.
@@ 3.2 Are all the layers necessary?
Figure 4: Left: Skipping layers N to 32-N for Llama2-7B, normalized for each benchmark, with median over all benchmarks. Right: Skipping layers N to 24-N for BERT, with unnormalized average.
@@ 3.3 Are middle layers all doing the same thing?
Figure 5: Replacing $M$ middle layers with the center layer (16 for Llama, 12 for BERT) for Llama2-7B (left, normalized benchmarks). and BERT (unnormalized average).
@@ 3.4 Does the layer order matter?
Figure 6: Left: Reversing $M$ middle layers for Llama2-7B, normalized across different Benchmarks. Right: Reversing layers for BERT-Large, unnormalized average.
Figure 7: Randomizing layer order for $M$ middle layers for Llama2-7B (left) and BERT (right). Each point is the average of 10 random seeds.
@@ 3.5 Can we run the layers in parallel?
Figure 8: Running $M$ layers (Layers $(T-M)/2$ to $(T-M)/2$) in parallel for Llama2-7B (left) and BERT (right)
@@ 3.6 Does the order matter for some tasks more than others?
In Appendix A.4 we show a specific example that indicates that errors for GSM8K may come from arithmetic errors in some cases.
@@ 3.7 Does looping help parallelized layers?
Figure 9: Running $M$ layers in parallel, looping 3 times for Llama2 (left) and BERT (right).
@@ 3.8 Which variants are least harmful?
Figure 11: Average benchmark scores for different variations for Llama2-7B (left) and BERT-large (right).
@@ 5 Discussion
In this paper, we examined several questions raised by the Layers as Painters analogy. Among our more interesting findings are: 1. There are three distinct classes of layers (with Middle being the largest). 2. The middle layers have some degree of uniformity (but not redundancy). And 3. Execution order matters more for math and reasoning tasks than semantic tasks.
@@ Appendix A Appendix
A.1 How does scaling affect hidden state similarities?
@@
A.2 Why is repeating a layer is so much worse than skipping it?
@@
Figure 13: Comparing average cosine similarity and variance between the hidden states of all layers of Llama2-7B in Skip (Left) and Middle Repeat (Right) methods, the start layer is 13 therefore Skip uses 27 layers, skipping 5.
A.3 Will internal looping improve over the base model?
Figure 14: Repeating layers N to 32-N for Llama2-7B 2 (left) and 3 (right) times.
No. We’re still below the full model baseline.
A.4 Example of Error for Parallelized Layers
@@
A.5 BERT benchmarks
@@
A.6 Results for Frozen BERT

### v3

@@ 1 Introduction
To answer these questions we perform a series of experiments on pretrained LLMs. These include experimenting with variations on the standard transformer execution strategy, and measuring the impact of these variations on the models’ performance across a variety of benchmarks for both decoder-only (Llama) and encoder-only (BERT) models. Note that our experiments never involve finetuning or otherwise adjusting the models’ parameters (with the caveat that the GLUE evaluation standard procedure includes a finetuning step for our BERT-Large model) 111 Our code is available at https://github.com/floatingbigcat/transformer-as-painter .
@@ 3.1 Do layers “speak the same language”?
Figure 3: Avg. cosine similarity between the hidden states of all 32 layers of Llama2-7B (top) and all 40 layers of Llama2-13B.
@@ 3.2 Are all the layers necessary?
In Appendix A.1, we analyze the layer skipping behavior across model sizes, revealing a surprisingly uniform pattern in the importance of middle layer partitions. Furthermore, in Appendix A.2, we demonstrate that fine-tuning can enhance performance when skipping fewer layers but becomes harmful when skipping too many.
Figure 4: Top: Skipping layers N to 32-N for Llama2-7B, normalized per benchmark (median). Bottom: Skipping layers N to 24-N for BERT, with unnormalized average.
@@ 3.3 Are middle layers all doing the same thing?
Figure 5: Replacing $M$ middle layers with the center layer (16 for Llama, 12 for BERT) for Llama2-7B (top, normalized benchmarks). and BERT (unnormalized average).
@@ 3.4 Does the layer order matter?
Figure 6: Top: Reversing $M$ middle layers for Llama2-7B, normalized across different Benchmarks. Bottom: Reversing layers for BERT-Large, unnormalized average.
Figure 7: Randomizing layer order for $M$ middle layers for Llama2-7B (top) and BERT (bottom). Each point is the average of 10 random seeds.
@@ 3.5 Can we run the layers in parallel?
Figure 8: Running $M$ layers (Layers $(T-M)/2$ to $(T-M)/2$) in parallel for Llama2-7B (top) and BERT (bottom)
@@ 3.6 Does the order matter for some tasks more than others?
In Appendix A.7 we show a specific example that indicates that errors for GSM8K may come from arithmetic errors.
@@ 3.7 Does looping help parallelized layers?
Figure 9: Running $M$ layers in parallel, looping 3 times for Llama2 (top) and BERT (bottom).
@@ 3.8 Which variants are least harmful?
Figure 11: Average benchmark scores for different variations for Llama2-7B (top) and BERT-large (bottom).
@@ 5 Discussion
In this paper, we examined several questions raised by the Layers as Painters analogy. Among our more interesting findings are: 1. There are three distinct classes of layers (with Middle being the largest). 2. The middle layers have some degree of uniformity (but not redundancy). And 3. Execution order matters more for math and reasoning tasks than semantic tasks. We welcome future theoretical analysis of layer behaviors in transformer architectures based on our empirical findings.
@@
Acknowledgements
We would like to thank Owen He, who came up with the painter analogy after seeing some of our early results. We would also like to thank Yujin Tang for providing valuable suggestions during the rebuttal process.
@@ Appendix A Appendix
A.1 Does skipping layers affect on the models changes across scales?
Emerging behavior of foundation models indicates different behaviors of models across sizes. An interesting question to ask is how does scaling affect the skipping intervention? We further launched the skip experiment for Llama2-13B and Llama2-70B to answer this question.
We unified the number of skipping layers across different model sizes into the percentage of skipped layers to get Figure 12. Surprisingly, all models showed similar trends in their retained performance, So we answer:
No, The affect remains consistent across scales, as middle layers partitions shows uniform importance regardless of model scale.
Figure 12: Skipping layers N to 32-N for Llama2-7B, 13B, and 70B.
A.2 How does fine-tuning affect the intervened model?
Fine-tuning is a powerful technique for improving model performance on specific tasks. And How does it affect the model with layer skipping? To understand this, we conduct experiments using Llama2-7B on the ARC challenge. We fully fine-tuned models with different numbers of skipped layers for $1$ epochs using a learning rate of $55e-$ 5 and batch size of $64$ on the ARC training set.
Figure 13 reveals that fine-tuning can indeed improve model robustness when fewer than $30\%$ of layers are skipped, with these models showing slower performance degradation compared to their frozen counterparts. We also noted that fine-tuning hurts performance when more than $30\%$ of layers are skipped. Therefore:
Fine-tuning benefits benign intervention but proves harmful for catastrophic intervention.
Figure 13: Comparison between directly skipping layers and fine-tuning the model after skipping layers on ARC challenge.
A.3 How does scaling affect hidden state similarities?
@@
A.4 Why is repeating a layer is so much worse than skipping it?
@@
Figure 15: Comparing average cosine similarity and variance between the hidden states of all layers of Llama2-7B in Skip (top) and Middle Repeat (bottom) methods, the start layer is 13 therefore Skip uses 27 layers, skipping 5.
@@
A.5 Do the results hold for Mistral and Pythia?
A.6 Will internal looping improve over the base model?
Figure 17: Repeating layers for Llama2-7B 2 and 3 times.
No. In Figure 17, we’re still below the full model baseline.
A.7 Example of Error for Parallelized Layers
A.8 BERT benchmarks
@@
A.9 Results for Frozen BERT